Finding the Dividend (LO1) Codner Corporation stock currently sells for $64 per share. The market requires a 10 percent return on the firm's stock. If the company maintains a constant 4.5 percent growth rate in dividends, what was the most recent dividend per share paid on the stock?
Answer #1

Required return=(D1/Current price)+Growth rate

0.1=(D1/64)+0.045

D1=(0.1-0.045)*64

=$3.52

Hence most recent dividend=D1*Present value of discounting factor(rate%,time period)

=3.52/1.045

=$3.37(Approx).